OpenStudy (anonymous):

personal income (in 1998 dollars) in one state increased approx linearly from $20,808 in 1998 to $22,395 in 2003. personal income (in 1998 dollars) in the other state increased approx linearly from $26,155 in 1998 to $26,508 in 2003. a) let f(t) be personal income (in 1998 dollars) in one state and g(t) be personal income (in 1998 dollars) in the other state, both in the year that is t years since 1998.Find the equation of f and g.

OpenStudy (anonymous):

Alright, so we have two points for f(t). The points are (0,20808) and (5,22395). The same with g(t); points (0,26155) and (0,26508). Do you understand this part?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

We know they are approximately linear. Therefore, we can right them as the first degree function f(x)=mx+b, where m is the slope, and b is the y-intercept--the y value when x=0. Slope can be found via the change in y over the change in x. In five years, how much money has accrued?

OpenStudy (anonymous):

i need both the f(t) and g(t) in an equation to do the substitution or elimination method

OpenStudy (anonymous):

and while figuring out the problem b is an outrageous number

OpenStudy (anonymous):

They're giving you outrageous numbers, so you should only expect outrageous numbers when you do math on them. ;D

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i did y-y1=m(x-x1)

OpenStudy (anonymous):

Alright, let me condense my hints. You're given two points for f(t) and g(t) each. Two points, (x1,y1), and (x2,y2). In this case, x1 is always x1=0, since the time t starts counting at the earliest year given. We know the form y=mx+b. m can be found a la (y2-y1)/(x2-x1), but remember x1 is 0. b can be found at f(0) or g(0), which we're already given; y1 for both cases. I'm sorry, but I'm not giving a direct answer. Knowing the method, and being careful with calculations, should be more than sufficient.

OpenStudy (anonymous):

it would be 22395-20808/5=317.4 f(t) 26508-26155/5= 70.6

OpenStudy (anonymous):

i got that part but when i plug it in to the equation it turns out to be this y-20808=317.4(x-0) right?

OpenStudy (anonymous):

Yeah, I think that works out right. Nicely done.

OpenStudy (anonymous):

when i multiply 317.4*20808 =6604459.2

OpenStudy (anonymous):

is that right for the equation?